# Feasible Space Monitoring for Multiple Control Barrier Functions with application to Large Scale Indoor Navigation

### v1

@@ VI-B Case Study 2: AWS Hospital
Next we evaluate our FS-CBF for navigating turtlebot3 robot in AWS Hospital gazebo environment shown in Fig. 1. We first generate a map of the environment without humans using ROS2 navigation stack. This map is used by global planner to plan paths, as well as by the CBF-QP controller to perform collision avoidance with static obstacles. At each time, the robot finds the closest occupied grid cell in the directions $\theta=\beta 30^{0},\beta\in\{1,2,..,12\}$, and formulates 12 CBF constraints correspondingly. The humans are moved in real-time using a modified version of social force model plugin [30] . The position and velocity of humans is extracted from gazebo and made available to the robot, which then adds a CBF constraint for each human. The proposed FS CBF-QP is compared with CBF-QP in (4). The collision avoidance is enforced using distance as the higher-order barrier function of order two which is known to work well in the presence of a single constraint only. For obstacles, we use linear class-$\mathcal{K}$ functions $\alpha_{1}(h)=2h,\alpha_{2}(h)=6h$ for first and second-order HOCBF derivative conditions. For humans, we simulate two scenarios: scenario 1 uses $\alpha_{1}(h)=2h,\alpha_{2}(h)=6h$ and scenario 2 uses $\alpha_{1}(h)=h,\alpha_{2}(h)=2h$. Scenario 1 controller is therefore more aggressive than that of Scenario 2. For each scenario, the experiment is run 15 times, and the performance statistics are reported in Table I. In each run, the controller start time is changed within an interval of $2$ seconds after humans start moving in the environment. The mean performance statistics are reported in Table I. The success rate is based on a robot rendered immobile after a hard collision in the gazebo. QP failures refer to the number of times the controller is infeasible over the run. Whenever the QP is infeasible, we implement the input from the previous time step. We see that FS-CBF-QP has better feasibility than CBF-QP. For the more aggressive scenario 1, the gains in feasibility are more pronounced compared to scenario 2. A video of the experiment is included in the supplementary material.
Method Success Rate QP Failures Success Rate QP Failures
Scenario 1 Scenario 2
CBF-QP 0.91 54.2 1.0 16.85
FS CBF-QP 1.0 34.5 1.0 15.42
TABLE I: CBF-QP vs proposed FS-CBF-QP on gazebo
VI-B1 Discussion of existing works
Note that several existing approaches such as [31, 32, 33] for indoor navigation with humans rely only on fast replanning through a low-level planner based on simplified dynamics for collision avoidance, and do not employ a safety-critical controller. A direct comparison with these approaches is therefore impractical. Our controller is thus meant to complement existing planning frameworks by providing them with a controller capable of not only enforcing collision avoidance but arbitrary state constraints for control-affine robot dynamics. Among CBF-based approaches, [34] uses CBF-MPC, an MPC controller with constraints defined by CBF derivative conditions, for navigation in crowd, and [35] designs a novel CBF for polytopic obstacles and uses it in the CBF-MPC framework for navigation in corridor environments. From an application perspective, [34] does not consider indoor environments and [35] does not consider dynamic moving agents. While MPC does help with maintaining feasibility of CBF constraints, it is unlike our method of monitoring the volume of feasible space. Furthermore, since no guarantees of recursive feasibility of MPC have been shown to exist in dynamic indoor environments, these controllers are still prone to infeasibility. Consideration of the volume of feasible space defined by CBF constraints may complement these nonlinear optimization-based MPC methods in improving feasibility but this analysis is left for future work and here we focus only on CBF-QP formulations.
@@ VII Conclusion
We propose a new barrier function that restricts the rate-of-change of the volume of feasible solution space of CBF-QP. We also provide empirical evaluation to corroborate our theoretical results, and to show effectiveness in improving feasibility and in reducing sensitivity to changes in other system modules such as nominal controller. Future work will involve performing experiments with robots and non-smooth control theoretic analysis of FS-CBF.

### v2

@@ VI-C Case Study 3: AWS Hospital
Next, we evaluate our FS-CBF for navigating turtlebot3 robot in AWS Hospital Gazebo environment shown in Fig. 1. We first generate a map of the environment without humans using ROS2 navigation stack. This map is used by global planner to plan paths, as well as by the CBF-QP controller to perform collision avoidance with static obstacles. At each time, the robot finds the closest occupied grid cell in the directions $\theta=\beta 30^{0},\beta\in\{1,2,..,12\}$, and formulates 12 CBF constraints correspondingly. The humans are moved in real-time using a modified version of social force model plugin [36]. Humans, however, are agnostic to robot’s motion and do not contribute to collision avoidance. The position and velocity of humans is extracted from Gazebo and made available to the robot, which then adds a CBF constraint for each human. The collision avoidance is enforced using distance as the higher-order barrier function of order two which is known to work well in the presence of a single constraint only. We use linear class-$\mathcal{K}$ functions $\alpha_{1}(h)=2h,\alpha_{2}(h)=6h$ for first and second-order HOCBF derivative conditions. The proposed FS CBF-QP is compared with CBF-QP in (5). We observe that CBF-QP leads to fewer collisions with humans by preventing the robot from getting too close to the humans in congested scenarios. A video of the experiment can be found at https://www.youtube.com/playlist?list=PLJxod9x5m_V2R5c4MLCCwwe_xliJgzOr2.
@@ VII Conclusion
We propose a new barrier function that restricts the rate-of-change of the volume of feasible solution space of CBF-QP. We also provide empirical evaluation to corroborate our theoretical results, and to show effectiveness in improving feasibility and in reducing sensitivity to changes in other system modules such as nominal controller. Future work will involve performing experiments with robots and non-smooth control theoretic analysis of FS-CBF. Finally, monitoring feasible space volume may also help alleviate feasibility issues in other optimization based controllers such as MPC. This analysis will also be performed in future work.